OpenStudy (seratul):

Can someone tell me if my steps are correct or not? "Solve the following system of equations: 3p-2q+2r=1 -2p+q-3r=2 3p-4q+3r=-7

OpenStudy (seratul):

Just reply to this if you can. :)

OpenStudy (3mar):

May I help?

OpenStudy (seratul):

Of course.

OpenStudy (seratul):

I wanted to answer using substitution.

OpenStudy (3mar):

Thank you. Can you share your works/steps, please?

OpenStudy (seratul):

So I made q=2+2p+3r I plugged it in for the first equation and got: 3p-2(2+2p+3r)+2r=1 3p-4-4p-6r+2r=1 -p=-5-4r p=-5-4r

OpenStudy (3mar):

you prefer sub. or ele.?

OpenStudy (seratul):

Sub.

OpenStudy (3mar):

Yes. I see!

OpenStudy (seratul):

Then for the third equation, I plugged in for q and got 3p-4(2+2p+3r)+3r=-7 -5p-9r=1

OpenStudy (seratul):

I plugged in for p for -5p-9r=1 and got -5(-5-4r)-9r=1 25+11r=1 r=(-24/11)

OpenStudy (seratul):

Good so far?

OpenStudy (seratul):

You still here?

OpenStudy (3mar):

p=-5-4r -5p-9r=1 \[\checkmark\]

OpenStudy (3mar):

I was checking one-by-one r=(-24/11) then sub. to get q and p...what are their values?

OpenStudy (seratul):

I got q=(32/11) and p=41/11

OpenStudy (3mar):

\[\Huge\color{greenyellow}\checkmark\] Excellent!

OpenStudy (3mar):

If there are any more questions, pleasure is mine to help!

563blackghost (563blackghost):

Nicely Done @Seratul

OpenStudy (3mar):

Thank you for the medal! You are Excellent!

OpenStudy (seratul):

My bad @3mar I had to go get something from the store. Thanks for the help!

OpenStudy (3mar):

No problem at all! You are welcome! Any Help... Any Time...